OpenStudy (anonymous):

18. What is an equation for the line that passes through points (8, –3) and (9, 10)? (1 point) (y + 3) = –13(x – 8) (y + 3) = 13(x – 8) (y – 3) = –13(x + 8) (y – 3) = 13(x + 8)

OpenStudy (anonymous):

is it.... C?

OpenStudy (anonymous):

hang on

OpenStudy (anonymous):

No it's the first one

OpenStudy (anonymous):

Let me expain:

OpenStudy (anonymous):

i mean explain:

OpenStudy (anonymous):

so to find the slope: 10--3 over 9-8 13/1=13 m=13

OpenStudy (anonymous):

then for point slope your formula is: y-y1=m(x-x1)

OpenStudy (anonymous):

So: y--3=13(x-8) and when you're subtracting a negative you're adding so: y+3=13(x-8)

OpenStudy (anonymous):

questions?

OpenStudy (anonymous):

oh so it was minus 8 not plus...derp!

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

So B is the right answer???

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

any more questions?

OpenStudy (anonymous):

thanks :D

OpenStudy (anonymous):

yep!